# StyleReiser: Stylizing Video With Reinforced Structure Guide

### v1

@@ 4 Results and Comparison
We implemented our approach in PyTorch Paszke et al. (2019) using the AdamW Loshchilov and Hutter (2017) optimizer with fixed learning rate $3\cdot 10^{-5}$. Since the batch size was set to $1$, we used instance normalization Ulyanov et al. (2016) instead of batch normalization. We set $\lambda_{k}=1.0$ and $\lambda_{v}=100.0$. The parameters for $\mathcal{L}_{\text{cSDS}}$ (3) were selected experimentally and differ between the presented sequences, $\lambda_{c}\in\{10^{-5},10^{-6}\}$, $t\in\{20,28\}$. For score sampling purposes, we adopt the default noise scheduler of ControlNet v1.1 Zhang et al. (2023b) and the UniPC scheduler Zhao et al. (2024) with the number of steps set to $30$. The parameters of the VGG network and ControlNet were fixed and lineart conditioning was used (if not stated otherwise). The training was performed on a single NVIDIA A100 GPU with $40$ GB of RAM for $4$ hours and the model with the lowest total loss has been selected to produce the stylized sequences. The results are presented in Figures 1, 3, 4, 5, and 6. See also our supplementary material for additional results. We compare them with the output of recent diffusion-based methods Ceylan et al. (2023); Chu et al. (2024); Geyer et al. (2024); Yang et al. (2023) and keyframe-based approaches Futschik et al. (2021); Jamriška et al. (2019); Texler et al. (2020b).
@@ 4.1 Comparison with diffusion-based methods
Since diffusion-based methods only support textual guidance, for each sequence presented, we provide a textual description $\mathbf{p}_{\text{desc}}$ that describes the content of the sequence. We also provide the edit description $\mathbf{p}_{\text{edit}}$ that guides the stylization process. The resulting text prompt is a concatenation of both $\mathbf{p}=\mathbf{p}_{\text{edit}}\mathbf{p}_{\text{desc}}$, e.g., $\mathbf{p}_{\text{edit}}=`$hyperrealistic detailed oil painting of```` and $\mathbf{p}_{\text{desc}}=`$an old man with a white beard````. The description $\mathbf{p}_{\text{desc}}$ is used as an inversion prompt in the cases of Ceylan et al. Ceylan et al. (2023) and Geyer et al. Geyer et al. (2024). The target video sequences, represented in all figures by a single frame $\mathbf{x}$, were stylized using the mentioned diffusion-based methods (to run them, we used the configuration scripts provided by the authors in their implementations). A single frame from each stylized sequence was selected as a keyframe and its style was propagated throughout the target sequence $\mathbf{x}\in\mathcal{X}$ using our method. From the results presented in Figures 3, 4, 5, and 6 it is apparent that our approach maintains structural details better than diffusion-based approaches. See also our supplementary video that demonstrates structural consistency across the entire sequence contrasting the flicker common in diffusion-based methods.
A key advantage of our approach in contrast to diffusion-based techniques is that it enables the user to incorporate custom edits by manually modifying the output of the diffusion-based method and use it as a newly stylized keyframe for training. This option could be highly beneficial, especially in cases when it is difficult to find an accurate text prompt that precisely expresses the desired artistic vision. In Fig. 7, we show the results of a custom edit of an image $\mathbf{x}_{r}$ stylized by the method of Yang et al. Yang et al. (2023). The stylized image $\mathbf{y}_{r}$ was generated with the text prompt “Galadriel, the royal Elf, silver-golden hair.” Custom edits were performed on the image $\mathbf{y}_{r}$ producing the image $\mathbf{y}_{e}$. Our method was then trained with the keyframe $(\mathbf{x}_{r},\mathbf{y}_{e})$, producing the stylization $\hat{\mathbf{y}}$.
@@ 4.2 Comparison with keyframe-based methods
For comparison with keyframe-based approaches (see Figures 8, 9, and 10) we used sequences generated by diffusion-based methods and then $4$ keyframes of the stylized sequence were selected to train the methods of Jamriška et al. Jamriška et al. (2019) and Texler et al. Texler et al. (2020b), and $1$ keyframe was selected to train the method of Futschik et al. Futschik et al. (2021) and our method. In all results presented, our approach consistently maintains the structural details of the target frame $\mathbf{x}$ while faithfully replicating important style features of the stylized keyframe. See also our supplementary video to compare the structural consistency across the entire sequence.
@@ 5 Experiments
In this section, we present a set of experiments that provide better insight into how our approach performs in various settings. We first examine its convergence rate (Sec. 5.1), then we present an ablation study on our loss components (Sec. 5.2). We investigate the influence of parameters $\lambda_{c}$ and $t$ (Sec. 5.3) and conditioning prior $\mathbf{c}$ (Sec. 5.4). Finally, we compare the results of training with $\mathcal{L}_{\text{cSDS}}$ and without it using only the lineart guidance function $c(\mathbf{x})$ (Sec. 5.5).
@@ 5.1 Convergence rate
In Fig. 12, we compare stylization results of our method (Figures 12a–c, blue color plot) with Futschik et al. Futschik et al. (2021) (Figures 12d–f, orange color plot) after $30$, $60$, and $90$ minutes of training (plots are averages over $15$ different sequences; the border of the opaque area is the standard deviation). The approach of Futschik et al. Futschik et al. (2021) emphasizes the preservation of the original style. However, this emphasis results in a gradual loss of fine details within the structure of the target sequence, as illustrated by the zoom-in insets in Fig. 12. Note how the eye in Fig. 12f is directly copied from the stylized keyframe $\mathbf{y}_{k}$. A key advantage of our approach is that we explicitly retain these fine structural details, thus maintaining fidelity to the target frames $\mathbf{x}$.
@@
Figure 12: Comparative analysis of stylization results over training duration. In each of the first three columns, a stylized target frame $\mathbf{x}$ is shown after $30$, $60$, and $90$ minutes of training. Top row (a), (b), (c) and blue color: our method, bottom row (d), (e), (f) and orange color: Futschik et al. Futschik et al. (2021). The method of Futschik et al. Futschik et al. (2021) prioritizes the preservation of the original style. However, over time, this leads to a gradual loss of fine details in the structure of the target sequence (cf. zoom-in insets). A key advantage of our method is that we explicitly strive to retain these details, thereby maintaining their fidelity. The plots of the total loss are averaged over $15$ different sequences, the border of the opaque area depicts the measured standard deviation.
@@ 5.2 Ablation study on our loss components
In this experiment, we performed an ablation study to analyze the impact of individual loss terms within the optimization of our model. We systematically set multiplication constant of each of the three loss terms $\lambda_{k}$, $\lambda_{v}$, and $\lambda_{c}$ (4) to zero. The results are presented in Fig. 13 for the model with the lowest total loss after training on a GPU for $4$ hours. We specifically selected an input frame $\mathbf{x}$ with a significant appearance change with respect to the style exemplar $\mathbf{y}$ to highlight the ability of our method to produce results in the style of the exemplar even when new, previously unseen structures appear in the input frame. The $t=28$ and $\lambda_{\text{cSDS}}=5\cdot 10^{-6}$.
@@
target frame $\mathbf{x}$
$\mathcal{L}_{\text{VGG}}+\mathcal{L}_{\text{cSDS}}$
$\mathcal{L}_{\text{key}}+\mathcal{L}_{\text{cSDS}}$
$\mathcal{L}_{\text{key}}+\mathcal{L}_{\text{VGG}}$
$\mathcal{L}_{\text{key}}+\mathcal{L}_{\text{VGG}}+\mathcal{L}_{\text{cSDS}}$
Figure 13: An ablation study demonstrating the importance of individual terms in our objective function (4). A neural network is trained to transfer the style from the stylized keyframe $\mathbf{y}$ to the target frame $\mathbf{x}$. Each of the three terms $\mathcal{L}_{\text{key}}$ (1), $\mathcal{L}_{\text{VGG}}$ (2), and $\mathcal{L}_{\text{cSDS}}$ (3) is left out in the training. Leaving out $\mathcal{L}_{\text{key}}$ (a) causes the style transfer to be less semantically meaningful (see, e.g., white nose), excluding $\mathcal{L}_{\text{VGG}}$ (b) leads to complete style washout, and leaving out $\mathcal{L}_{\text{cSDS}}$ (c) results in poor replication of target frame structures. Only the combination of all three terms yield satisfactory results (d).
Excluding $\mathcal{L}_{\text{key}}$ from the loss in Fig. 13a results in an output that preserves structure of the target frame $\mathbf{x}$, however, the transfer is not semantically meaningful, i.e., colors from the stylized keyframe $\mathbf{y}$ are placed to improper locations in the target frame $\mathbf{x}$ (cf. white spots located on the nose and forehead).
When the term $\mathcal{L}_{\text{VGG}}$ is omitted in Fig. 13b, we get a result with an overall structure of the target frame $\mathbf{x}$ but with corrupted style details. This is caused by the fact that the term $\mathcal{L}_{\text{cSDS}}$ enforces the structure of the target frame $\mathbf{x}$ but not details of the stylized keyframe $\mathbf{y}$ that are maintained by the term $\mathcal{L}_{\text{VGG}}$.
When $\mathcal{L}_{\text{cSDS}}$ is omitted in Fig. 13c, the term $\mathcal{L}_{\text{VGG}}$ enforces the presence of the style details in the keyframe image $\mathbf{y}$, and the term $\mathcal{L}_{\text{key}}$ keeps the style transfer to be semantically meaningful. However, when previously unseen structures appear, such as new shapes of eyes, they are poorly reconstructed since the loss without term $\mathcal{L}_{\text{cSDS}}$ forces the network to reproduce only structures appearing in the stylized keyframe $\mathbf{y}$.
@@
5.3 The influence of parameters $\lambda_{c}$ and $t$
In this experiment, we manipulate the parameter $t$ in $\mathcal{L}_{\text{cSDS}}$ (3) and the loss weight $\lambda_{c}$ (4). We trained the network for $4$ hours on a GPU and selected the network with the lowest total loss.
@@ 5.3 The influence of parameters $\lambda_{c}$ and $t$
$\lambda_{c}=5\cdot 10^{-5}$
$\lambda_{c}=1\cdot 10^{-5}$
$\lambda_{c}=5\cdot 10^{-6}$
$\lambda_{c}=1\cdot 10^{-6}$
Figure 14: Reinforcement of input video structure as function of parameters $\lambda_{c}$ (4) and $t$ (3). The stronger $\lambda_{c}$, the more pronounced the structure of $\mathbf{x}$ is in the result. Note that the ControlNet used by the structural guidance term $\mathcal{L}_{\text{cSDS}}$ (3) uses a $30$ step denoising schedule. At $t=0$, the guidance input is pure Gaussian noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and at $t=29$, it contains minimal noise.
The results in Fig. 14 show that the stronger $\lambda_{c}$, the more pronounced the structure of $\mathbf{x}$ is in the result. Since the value of $\lambda_{c}$ modifies the strength of $\mathcal{L}_{\text{cSDS}}$ in the total loss term $\mathcal{L}$ (4), setting it to a higher value results in a stronger pronunciation of the input video structures in the output of our method. Another interesting dimension of structure reinforcement control strength is the parameter $t$. The ControlNet we utilize for the purpose of structure guidance uses a $30$ step denoising schedule. Every time the loss term $\mathcal{L}_{\text{cSDS}}$ is evaluated, a convex combination $\hat{\mathbf{y}}_{i,t}$ (3) of the output of our method and a Gaussian noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is performed. Then, the pre-trained ControlNet estimates the original noise $\epsilon$, and the residual between the estimated noise and $\epsilon$ is propagated through our method. This means that when $t=0$, ControlNet gets a pure Gaussian noise and when $t=29$, it gets the output of our network with only a little Gaussian noise present. According to this, we see that the lower the value of $t$, the stronger the focus is on the high-level structures compared to the higher values of $t$.
@@ 5.4 The influence of image conditioning prior $\mathbf{c}$
In our method, we use a ControlNet diffusion network Zhang et al. (2023b) to enforce the appearance of the structural elements of the target frame $\mathbf{x}$ in the stylized output of our network. We hypothesize that through sampling from the ControlNet network, we take advantage of both the structural conditioning provided by the guidance image $\mathbf{c}$ and the tendency of the network to guide the denoised samples toward the learned manifold of real images.
In this experiment, we investigate the suitability of four versions of pre-trained ControlNet models for stylization structural guidance. Each pre-trained ControlNet uses a different image guidance function $c$: (i) “lineart detection” neural model, (ii) “depth estimation” neural model, (iii) Canny edge detector Canny (1986), and (iv) “soft edge estimation“ neural model.
The results for $t=16$ and $t=28$ are presented in Fig. 15. Both depth and soft edge conditioning produce stylization that lacks details of the target frame $\mathbf{x}$. This appears to be caused by a very coarse structure of the two types of guidance, as seen in the guidance image $\mathbf{c}$. The Canny edge detector adds more signal to the ControlNet guidance, and the image guidance with the largest amount of signal is the lineart. Note that in applications where a fast training time is paramount, the quality of the lineart results might be traded for $\approx 10$ times faster Canny edge detector guidance. The depth guidance is in this sense impractical as it takes $\approx 10$ times more computational time than the lineart guidance, with the soft edge guidance somewhere in between the two.
In this experiment, we kept the $\lambda_{c}=5\cdot 10^{-6}$ and trained on a single NVIDIA A100 GPU for $4$ hours. We selected the model with the lowest total loss.
@@
Figure 15: Structure reinforcement as function of the image guidance type $\mathbf{c}$ and the number of diffusion time steps $t$. The best results are obtained using the lineart guidance that provides the ControlNet in $\mathcal{L}_{\text{cSDS}}$ (3) with the most detailed structural image and $t=28$.

### v2

@@ 4 Results and Comparison
The results are presented in Figures 1, 3, 4, 5, and 6. See also our supplementary material for additional results. We compare them with the output of recent diffusion-based methods Ceylan et al. (2023); Chu et al. (2024); Geyer et al. (2024); Yang et al. (2023) and keyframe-based approaches Futschik et al. (2021); Jamriška et al. (2019); Texler et al. (2020b).
4.1 Perceptual study
To qualitatively evaluate our approach, we conducted a perceptual study comparing our method’s outputs with those from four state-of-the-art diffusion-based techniques (Ceylan et al. (2023); Chu et al. (2024); Geyer et al. (2024); Yang et al. (2023)), and three state-of-the-art exemplar-based methods ( Futschik et al. (2021); Jamriška et al. (2019); Texler et al. (2020b)). The study assessed how well each method reproduced the artistic style, preserved structural content, and maintained temporal consistency of the input video. We collected data from $54$ participants through an online survey, where participants were presented with randomized two-alternative forced-choice (2AFC) comparisons. Each participant completed $36$ questions, selecting which anonymized stylization better reproduced style (12 questions), preserved content (12 questions), and maitained temporal consistency (12 questions). In each comparison, an output from our method was paired with one from another method using the same input data.
The preference scores for our method versus others are presented in Fig. 11. In the figure, hollow squares represent preferences regarding temporal consistency, while hollow triangles indicate preferences concerning structural elements. Our method significantly outperforms previous works in reproducing input structures, even though it reproduces the exemplar’s style slightly less accurately – which is expected due to our focus on structural preservation. Moreover, our method excels in maintaining the temporal consistency of the stylized output compared to previous approaches.
4.2 Quantitative evaluation
To quantitatively evaluate of our method’s ability to preserve structural elements from input videos, we conducted a comprehensive comparison across published video stylization techniques. We calculated the averages and standard deviations between corresponding input and stylized frames across all available video sequences, using three metrics: SSIM, LPIPS zhang_unreasonable_2018, and F LIP andersson_flip_2020. The style exemplars were sourced from Geyer et al. (2024), Yang et al. (2023), Ceylan et al. (2023), Chu et al. (2024), and a combination including artist-created stylizations.
Since the published diffusion-based methods cannot perform exemplar-based stylization, we trained our method using a single style exemplar selected from the output of each diffusion method. We then compared the results of our method with those of the diffusion-based methods, as shown in the first four groups of results in Table 1. In the last group, we compare our method with the exemplar-based stylization approaches of Jamriška et al. Jamriška et al. (2019), Futschik et al. Futschik et al. (2021), and Texler et al. Texler et al. (2020b), trained on a set of diffusion-based exemplars and artist-created exemplars. Our method demonstrates superior performance across all sequence groups and metrics, effectively maintaining structural fidelity across diverse style sources.
@@ 4.3 Comparison with diffusion-based methods
Since diffusion-based methods only support textual guidance and cannot perform keyframe-based stylization directly, we adopt a two-step approach for each video sequence. First, we generate a stylized version using a combined text prompt $\mathbf{p}=\mathbf{p}_{\text{edit}}\mathbf{p}_{\text{desc}}$, where $\mathbf{p}_{\text{edit}}$ (e.g., “hyperrealistic detailed oil painting of”) defines the desired style, and $\mathbf{p}_{\text{desc}}$ (e.g., ”an old man with a white beard“) describes the content. For instance, $\mathbf{p}=$ “hyperrealistic detailed oil painting of an old man with a white beard”. From the resulting stylized sequence, we then select one frame as a keyframe and propagate this keyframe’s style throughout the entire target sequence $\mathbf{y}\in\mathcal{Y}$ using our proposed keyframe-based stylization method. As a result, each diffusion-based method is presented with a unique stylization. Note that in the methods proposed by Ceylan et al. Ceylan et al. (2023) and Geyer et al. Geyer et al. (2024), the context description $\mathbf{p}_{\text{desc}}$ serves as an inversion prompt.
From the results presented in Figures 3, 4, 5, and 6 it is apparent that our approach maintains structural details better than diffusion-based approaches. See also our supplementary video that demonstrates structural consistency across the entire sequence contrasting the flicker common in diffusion-based methods.
A key advantage of our approach in contrast to diffusion-based techniques is that it enables the user to incorporate custom edits by manually modifying the output of the diffusion-based method and use it as a newly stylized keyframe for training. This option is beneficial especially in cases when it is difficult to find an accurate text prompt that precisely expresses the desired artistic vision. In Fig. 7, we show the results of a custom edit of an image $\hat{\mathbf{x}}_{r}$ stylized by the method of Yang et al. Yang et al. (2023) with the text prompt “Galadriel, the royal Elf, silver-golden hair.” This image was edited by an artist producing the image $\hat{\mathbf{x}}_{e}$. Our method was then trained with the keyframe $(\mathbf{x}_{r},\hat{\mathbf{x}}_{e})$, rendering the stylization $\hat{\mathbf{y}}$.
@@ 4.4 Comparison with keyframe-based methods
To compare our method with other keyframe-based approaches (see Figures 8, 9, and 10), we used sequences generated by diffusion-based methods. From each stylized sequence, we selected keyframe to train different methods: four keyframes were chosen to train the methods of Jamriška et al. Jamriška et al. (2019) and Texler et al. Texler et al. (2020b), while one keyframe was used to train the method of Futschik et al. Futschik et al. (2021) and our own method. In all the results presented, our approach consistently maintains the structural details of the target frame $\mathbf{y}$ while faithfully replicating important style features of the stylized keyframe. Please refer to our supplementary video to compare the structural consistency across the entire sequence.
@@ 5 Experiments
In this section, we present a set of experiments that provide better insight into how our approach performs in various settings. We first examine its convergence rate (Sec. 5.1), then we present an ablation study on our loss components (Sec. 5.2). We investigate the influence of parameters $\lambda_{s}$ and $t$ (Sec. 5.3) and conditioning prior $\mathbf{c}$ (Sec. 5.4). Finally, we compare the results of training with $\mathcal{L}_{\text{structure}}$ and without it using only the line art guidance function $c(\mathbf{y})$ (Sec. 5.5).
@@ 5.1 Convergence rate
In Fig. 13, we compare stylization results of our method (Figures 13a–c, blue color plot) with Futschik et al. Futschik et al. (2021) (Figures 13d–f, orange color plot) after $30$, $60$, and $90$ minutes of training (plots are averages over $15$ different sequences; the border of the opaque area is the standard deviation). The approach of Futschik et al. Futschik et al. (2021) emphasizes the preservation of the original style. However, this emphasis results in a gradual loss of fine details within the structure of the target sequence, as illustrated by the zoom-in insets in Fig. 13. Note how the eye in Fig. 13f is directly copied from the stylized keyframe $\mathbf{y}_{k}$. A key advantage of our approach is that we explicitly retain these fine structural details, thus maintaining fidelity to the target frames $\mathbf{y}$.
@@
Figure 13: Comparative analysis of stylization results over training duration. In each of the first three columns, a stylized target frame $\mathbf{y}$ is shown after $30$, $60$, and $90$ minutes of training. Top row (a), (b), (c) and blue color: our method, bottom row (d), (e), (f) and orange color: Futschik et al. Futschik et al. (2021). The method of Futschik et al. Futschik et al. (2021) prioritizes the preservation of the original style. However, over time, this leads to a gradual loss of fine details in the structure of the target sequence (cf. zoom-in insets). A key advantage of our method is that we explicitly strive to retain these details, thereby maintaining their fidelity. The plots of the total loss are averaged over $15$ different sequences, the border of the opaque area depicts the measured standard deviation.
@@ 5.2 Ablation study on our loss components
In this experiment, we performed an ablation study to analyze the impact of individual loss terms within the optimization of our model. We systematically set multiplication constant of each of the three loss terms $\lambda_{k}$, $\lambda_{v}$, and $\lambda_{s}$ (1) to zero. The results are presented in Fig. 14 for the model with the lowest total loss after training on a GPU for $4$ hours. We specifically selected an input frame $\mathbf{x}$ with a significant appearance change with respect to the style exemplar $\mathbf{y}$ to highlight the ability of our method to produce results in the style of the exemplar even when new, previously unseen structures appear in the input frame. The $t=28$ and $\lambda_{\text{structure}}=5\cdot 10^{-6}$.
@@
target frame $\mathbf{y}$
$\mathcal{L}_{\text{style}}+\mathcal{L}_{\text{structure}}$
$\mathcal{L}_{\text{key}}+\mathcal{L}_{\text{structure}}$
$\mathcal{L}_{\text{key}}+\mathcal{L}_{\text{style}}$
$\mathcal{L}_{\text{key}}+\mathcal{L}_{\text{style}}+\mathcal{L}_{\text{structure}}$
Figure 14: An ablation study demonstrating the importance of individual terms in our objective function (1). A neural network is trained to transfer the style from the stylized keyframe $\mathbf{y}$ to the target frame $\mathbf{y}$. Each of the three terms $\mathcal{L}_{\text{key}}$ (2), $\mathcal{L}_{\text{style}}$ (3), and $\mathcal{L}_{\text{structure}}$ (4) is left out in the training. Leaving out $\mathcal{L}_{\text{key}}$ (a) causes the style transfer to be less semantically meaningful (see, e.g., white nose), excluding $\mathcal{L}_{\text{style}}$ (b) leads to complete style washout, and leaving out $\mathcal{L}_{\text{structure}}$ (c) results in poor replication of target frame structures. Only the combination of all three terms yield satisfactory results (d).
Excluding $\mathcal{L}_{\text{key}}$ from the loss in Fig. 14a results in an output that preserves structure of the target frame $\mathbf{y}$, however, the transfer is not semantically meaningful, i.e., colors from the stylized keyframe $\mathbf{y}$ are placed to improper locations in the target frame $\mathbf{y}$ (cf. white spots located on the nose and forehead).
When the term $\mathcal{L}_{\text{style}}$ is omitted in Fig. 14b, we get a result with an overall structure of the target frame $\mathbf{y}$ but with corrupted style details. This is caused by the fact that the term $\mathcal{L}_{\text{structure}}$ enforces the structure of the target frame $\mathbf{y}$ but not details of the stylized keyframe $\mathbf{y}$ that are maintained by the term $\mathcal{L}_{\text{style}}$.
When $\mathcal{L}_{\text{structure}}$ is omitted in Fig. 14c, the term $\mathcal{L}_{\text{style}}$ enforces the presence of the style details in the keyframe image $\mathbf{y}$, and the term $\mathcal{L}_{\text{key}}$ keeps the style transfer to be semantically meaningful. However, when previously unseen structures appear, such as new shapes of eyes, they are poorly reconstructed since the loss without term $\mathcal{L}_{\text{structure}}$ forces the network to reproduce only structures appearing in the stylized keyframe $\mathbf{y}$.
@@
5.3 The influence of parameters $\lambda_{s}$ and $t$
In this experiment, we manipulate the parameter $t$ in $\mathcal{L}_{\text{structure}}$ (4) and the loss weight $\lambda_{s}$ (1). We trained the network for $4$ hours on a GPU and selected the network with the lowest total loss.
@@ 5.3 The influence of parameters $\lambda_{s}$ and $t$
$\lambda_{s}=5\cdot 10^{-5}$
$\lambda_{s}=1\cdot 10^{-5}$
$\lambda_{s}=5\cdot 10^{-6}$
$\lambda_{s}=1\cdot 10^{-6}$
Figure 15: Reinforcement of input video structure as function of parameters $\lambda_{s}$ (1) and $t$ (4). The stronger $\lambda_{s}$, the more pronounced the structure of $\mathbf{x}$ is in the result. Note that the ControlNet used by the structural guidance term $\mathcal{L}_{\text{structure}}$ (4) uses a $30$ step denoising schedule. At $t=0$, the guidance input is pure Gaussian noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and at $t=29$, it contains minimal noise.
The results in Fig. 15 show that the stronger $\lambda_{s}$, the more pronounced the structure of $\mathbf{x}$ is in the result. Since the value of $\lambda_{s}$ modifies the strength of $\mathcal{L}_{\text{structure}}$ in the total loss term $\mathcal{L}$ (1), setting it to a higher value results in a stronger pronunciation of the input video structures in the output of our method. Another interesting dimension of structure reinforcement control strength is the parameter $t$. The ControlNet we utilize for the purpose of structure guidance uses a $30$ step denoising schedule. Every time the loss term $\mathcal{L}_{\text{structure}}$ is evaluated, a convex combination $\hat{\mathbf{y}}_{i,t}$ (4) of the output of our method and a Gaussian noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is performed. Then, the pre-trained ControlNet estimates the original noise $\epsilon$, and the residual between the estimated noise and $\epsilon$ is propagated through our method. This means that when $t=0$, ControlNet gets a pure Gaussian noise and when $t=29$, it gets the output of our network with only a little Gaussian noise present. According to this, we see that the lower the value of $t$, the stronger the focus is on the high-level structures compared to the higher values of $t$.
@@ 5.4 The influence of image conditioning prior $\mathbf{c}$
In our method, we use a ControlNet diffusion network Zhang et al. (2023b) to ensure that the structural elements of the target frame $\mathbf{y}$ are accurately reflected in the stylized output. We hypothesize that by sampling from the ControlNet network, we leverage both the structural conditioning provided by the guidance image $\mathbf{c}$ and the network’s inherent ability to guide denoised samples toward the learned manifold of real images.
In this experiment, we evaluate the effectiveness of four pre-trained ControlNet models for providing structural guidance in stylization. Each ControlNet uses a different image guidance function $c$: (i) “line art detection” neural model, (ii) “depth estimation” neural model, (iii) Canny edge detector Canny (1986), and (iv) “soft edge estimation“ neural model.
The results for $t=16$ and $t=28$ are presented in Fig. 16. Both depth and soft edge conditioning result in stylization that lacks the finer details of the target frame $\mathbf{y}$, likely due to the coarse structure of their respective guidance images $\mathbf{c}$. The Canny edge detector provides more detailed signal to the ControlNet, while line art guidance offers the most precise structural information. For applications requiring rapid training, the quality of line art can be sacrificed for the approximately $10$-fold faster Canny edge guidance. In contrast, depth guidance is less practical, taking about $10$ times longer than line art, with soft edge guidance falling between the two in terms of computational cost.
In this experiment, we kept $\lambda_{s}=5\cdot 10^{-6}$ and trained on a single NVIDIA A100 GPU for $4$ hours. We selected the model with the lowest total loss.
@@
Figure 16: Structure reinforcement as function of image guidance $\mathbf{c}$ and diffusion time steps $t$. The best results are achieved with the line art guidance within ControlNet’s $\mathcal{L}_{\text{structure}}$ (4) at $t=28$.